OpenStudy (mathmath333):

question(logical reasoning) ABC paints have came up with brilliant marketing...

OpenStudy (knov):

question?

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated & & 1\ jan & emulsion \\ \hline la-lina & & & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & & & \\ \hline \end{array} \end{align}}\)

OpenStudy (mathmath333):

@knov @reemii

OpenStudy (reemii):

The last information (vi) is not necessary because we already know Phoenix will launch 'Duplex'. Still thinking …

OpenStudy (reemii):

DD launches on January 1st because: (i) Launch is in February (ii) Launch is in March (ii) "After UK, < June 14" => not January. (iii) "before SouthAfrica, not January1st" => not January (v) "Launch is not in January"

OpenStudy (mathmath333):

ok so \(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated & & 1\ jan & emulsion \\ \hline la-lina & & & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & &1\ jan & \\ \hline \end{array} \end{align}}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated & & & emulsion \\ \hline la-lina & & & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & &1\ jan & \\ \hline \end{array} \end{align}}\)

OpenStudy (reemii):

Now, look at the launch date of Nina. (iii).It must be before the launch date in SouthAfrica. (iv) DD launches in Australia or SouthAfrica. => SouthAfrica is impossible: Nina can't launch before January1st. => DD launches in Australia. Second thing. (iii) launches before SouthAfrica -> it does not launch in SouthAfrica. What can it be then? => find country for RR.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated & & & emulsion \\ \hline la-lina & & & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & australia &1\ jan & \\ \hline \end{array} \end{align}}\)

OpenStudy (reemii):

Yes. the two countries that remain are Brazil and SouthAfrica. But Nina launches before SouthAfrica.. -> Nina is Brazil and RR is SouthAfrica.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated & & & emulsion \\ \hline la-lina & & & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & australia &1\ jan & \\ \hline \end{array} \end{align}}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated &south-africa & & emulsion \\ \hline la-lina & brazil & & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & australia &1\ jan & \\ \hline \end{array} \end{align}}\)

OpenStudy (mathmath333):

what about the next two que

OpenStudy (reemii):

Now study the months of launch. Jan, Feb, Mar already taken. Use the other information to find the dates. For example, Nina: before SouthAfrica, after January. => only April and May remain. Look at the the info for RR. What do you find?

OpenStudy (mathmath333):

rr is not on 14 june

OpenStudy (reemii):

so it's April or May. Nina also: April or May. But Nina launches before SouthAfrica( = RR)

OpenStudy (mathmath333):

rr=may ?

OpenStudy (reemii):

I think so. And Nina -> April

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|} \hline sonya & usa & 5\ feb & \\ \hline maxwell & uk & 23\ mar & \\ \hline r-rated &south-africa & 22\ may & emulsion \\ \hline la-lina & brazil & 29\ april & velvet \\ \hline phoenix & iv-coast & & duplex \\ \hline d-damels & australia &1\ jan & \\ \hline \end{array} \end{align}}\)

OpenStudy (reemii):

Yes. You can now answer the two last questions.

OpenStudy (mathmath333):

brazil -29

OpenStudy (mathmath333):

14 jun-duplex

OpenStudy (mathmath333):

is it correct

OpenStudy (reemii):

I would say that too. I hope it's correct. :)

OpenStudy (reemii):

You don't need to fill the table, right?

OpenStudy (mathmath333):

book answer ^

OpenStudy (mathmath333):

no i only nneed answer of the que

OpenStudy (mathmath333):

just to confirm that ur answers were correct :)

OpenStudy (mathmath333):

thnx

OpenStudy (reemii):

yw